OpenStudy (anonymous):

Lesson 7: Geometry B Unit 1: Similarity? If you help me, I'll help you! :)

OpenStudy (anonymous):

what do you need help with about similarity?

OpenStudy (anonymous):

The answer would be "d".

OpenStudy (anonymous):

are you in 10th grade?

OpenStudy (anonymous):

I have Geom B, why?

OpenStudy (anonymous):

im in the same thing and I have my exams tomorrow and need help. I do online school so its hard to get the one on one attention I need sometimes.

OpenStudy (anonymous):

I do online too.

OpenStudy (anonymous):

Do you go to Northstar?

OpenStudy (anonymous):

Nope :/ Connexus

OpenStudy (anonymous):

Well I do connexus but the school is called northstar.

OpenStudy (anonymous):

Hmm. That's cool! :) lol. Message me and and we can see if we can help eachother

OpenStudy (anonymous):

ok do you do your proficiency exams tomorrow too?

OpenStudy (anonymous):

uuumm. No.lol. I'm just in my unit 1 test right now and I'm stuck.

OpenStudy (anonymous):

hey I've been there I barely started on jan. 27 and now I have my proficiency exam. :0

OpenStudy (anonymous):

Oh, man! :( I wasn't able to do stuff for a while so I'm a little behind. It all kind of sucks.lol

OpenStudy (anonymous):

same here I wasn't in school for a little while and ended up being behind but everyone says I caught up pretty quick.

OpenStudy (anonymous):

The classes (except this one) are pretty easy. I have a one year old running around so it's a little hard for me to catch up! But I have nothing below a B, so I guess I'm doing good.lol

OpenStudy (anonymous):

im do great in English and foreign languages, okay in history and science, but math is the one thing I am confuzzled in. :)

OpenStudy (anonymous):

I haaatttee Math. I'm pretty good with Algebra but it takes a while to set in!

OpenStudy (anonymous):

Exactly! By the way did you get that question right?

OpenStudy (anonymous):

I don't know, I haven't finished the test. :(

OpenStudy (phi):

there are 3 right triangles in your picture, and *all three* are similar to each other. The biggest triangle is the outer triangle. the other 2 are inside the big one. the smallest triangle with sides x,y,8 has short leg to long leg x:8 the middle size triangle has sizes 8,15,17. short leg to long leg is 8:15 we can write a ratio \[ \frac{x}{8}= \frac{8}{15} \] if you multiply both sides by 8 \[ \frac{x}{8}\cdot 8 = \frac{8}{15} \cdot 8 \] you get \[ x = \frac{64}{15} \]